You will use your Minitab graphs created in Lesson 1 to answer the following questions.

 

Frequency Histogram Use your frequency histogram to answer the following questions. 

a) How many women were in the sample? 

b) What glucose level (bar) is at the center of the data? Identify the class limits. 

c) what is the shape of the distribution? 

d) How many glucose levels are in the modal (largest) class? 

e) How many glucose levels are in the smallest class? 

f) How many glucose levels are at most 65 mg/100 ml? 

g) How many glucose levels are between 85 and 95 mg/100 ml? 

h) How many glucose levels are at least 95 mg/100 ml? 


Relative Frequency Histogram 

Use your relative frequency histogram to answer the following questions. Round percentages to 2 decimals.

a) what is the difference between the frequency histogram and the relative frequency histogram?

b) What is the sum of all the bar percentages? 

c) What percent of glucose levels are in the modal (largest) class? 

d) What percent of glucose levels are in the smallest class? 

e) What percent of glucose levels are between 65 and 75 mg/100 ml? 

f) What percent of glucose levels are at most 75 mg/100 mi?

g What percent of glucose levels are at least 95 mg/100 ml?

Histogram of Glucose Level Random Blood Glocuse Level Sample 14 14 12 12 10 10- 戸: 6 2- 0- 50 70 80 90 100 110 Glucose LevelHistogram of Glucose Level Random Blood Glocuse Level Sample 30- 27 25 19 20 12 10- 45 65 75 85 95 105 115 Glucose Level2 Histogram of Glucose Level Random Blood Glocuse Level Sample 40 38 5714 300 27.1429 17.1429 10 5.71429 5.71429 4.28571 142857 45 65 75 85 95 105 115 Glucose LevelYou will use your Minitab graphs created in Lesson 1 to answer the following questions Frequency Histogram Use your frequency histogram to answer the following questions. a) How many women were in the sample? b) What glucose level (bar) is at the center of the data? Identify the class limits. o) what is the shape of the distribution? d) How many glucose levels are in the modal (largest) class? e) How many glucose levels are in the smallest class? How many glucose levels are at most 65 mg/100 ml? g) How many glucose levels are between 85 and 95 mg/100 ml? h) How many glucose levels are at least 95 mg/100 ml? Relative Frequency Histogram Use your relative frequency histogram to answer the following questions. Round percentages to 2 decimals a) what is the diference between the frequency histogram and the re b) What is the sum of all the bar percentages? c) What percent of glucose levels are in the modal (largest) class? d) What percent of glucose levels are in the smallest class? e) What percent of glucose levels are between 65 and 75 mg/100 ml? What percent of glucose levels are at most 75 mg/100 mi?g What percent of glucose levels are at least 95 mg/100 ml?
Answer #1

Frequency Histogram

a) How many women were in the sample? Solution: N= (=3+0+0+2+6+10+11+14+12+3+5+3+0+1)= 70 or N=3+4+19+27+12+4+1=70

b) What glucose level(bar) is at the center of the data? Identify the class limits. Solution: Glucose level= 80 [ (75+85)/2 ]

Class limits: 75-85 ( From graph 1 or 2 )

c) What is the shape of the distribution? Solution: Normal Distribution. ( Bell-shaped curve)

d) How many glucose levels are there in the modal class(largest class)? Solution: Number of women with glucose levels in Modal class i.e the class which has the highest frequency i.e 27 Class: 75-85 ( ( From graph 2 )

e) How many glucose levels are there in the smallest class? Solution : 1 Class : 105-115 ( From graph 2 )

f) How many glucose levels are at most 65mg/100ml? Solution : Glucose levels are at most 65mg/100ml mean Women havingGlucose levels less than or equal to 65mg =3+4 =7 ( From graph 2 )

g) How many glucose levels are between 85 and 95 mg/100ml? Solution: 12 ( From graph 2 )

h) How many glucose levels are at least 95mg/100ml? Solution: Women with Glucose levels greater than or equal to 95mg/100ml = 4+1=5 ( From graph 2 )

Relative Frequency Histogram :

a) Difference between Frequency Histogram and Relative Frequency Histogram

Frequency Histogram Relative Frequency Histogram
Uses simple frequency or number of observations Uses relative frequency or proportional frequency i.e Frequency of each class interval / Total Frequency
Horizontal axis: X-axis: Count of observations Horizontal axis: X-axis: Percentage of observations

b) What is the sum of all the bar percentages? Solution: Sum is always 100%

c) How many glucose levels are there in the modal class(largest class)? Solution: Number of women with glucose levels in Modal class i.e the class which has the highest frequency i.e 27 Class: 75-85 ( ( From graph 2 )

d) How many glucose levels are there in the smallest class? Solution : 1 Class : 105-115 ( From graph 2 )

e) How many glucose levels are at most 65mg/100ml? Solution: Glucose levels are at most 65mg/100ml mean Women having glucose levels less than or equal to 65mg =3+4 =7 ( From graph 2 )

f) How many glucose levels are between 85 and 95 mg/100ml? Solution: 12 ( From graph 2 )

g) How many glucose levels are at least 95mg/100ml? Solution: Women with Glucose levels greater than or equal to 95mg/100ml = 4+1=5 (From graph 2 )